How can i calculate (d2y/dx2) of any function y=f(x)

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chandrakesh

 meaning of d2y/dxis

d2y/dx= d/dx (dy/dx)

 

for example we have a function  y =  x4 + 20x2

now we need to find out d2y/dx

first find out dy/dx   

    so dy/dx = 4x3 + 40x

now for d2y/dxdifferentiate the dy/dx again

     so  d2y/dx2 = d/dx ( 4x3 + 40x)

                     = 12x2 +40

 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin